OpenStudy (anonymous):

PLEASE HELP Describe how to transform the quantity of the sixth root of x to the fifth power, to the seventh power into an expression with a rational exponent.

OpenStudy (aum):

\[\large \left(\sqrt[6]{x^5}\right)^7 = \left((x^5)^{1/6}\right)^7 = \left(x^{5/6}\right)^7 = x^{35/6} \]